OpenStudy (anonymous):

please help log8 x^3=-2 log(x-3)+logx=log4

OpenStudy (asnaseer):

do you know the various log rules?

OpenStudy (asnaseer):

if not, you can find them here: http://www.purplemath.com/modules/logrules.htm they should help you solve this.

OpenStudy (asnaseer):

please ask if you need any clarification

OpenStudy (anonymous):

still cant understand those rules sorry

OpenStudy (asnaseer):

which part confuses you?

OpenStudy (anonymous):

the one to solve the first problem

OpenStudy (asnaseer):

ok, I tihkn your first problem is:\[\log_8(x^3)=-2\]correct?

OpenStudy (asnaseer):

*think

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

right, lets break this up into parts. firstly, if you have:\[\log_a(x^b)\]then this can be simplified to:\[b*log_a(x)\]

OpenStudy (asnaseer):

so we simplify you 1st equation to:\[3\log_8(x)=-2\]

OpenStudy (anonymous):

I am solving for x

OpenStudy (asnaseer):

next we can divide both sides by 3 to get:\[\log_8(x)=\frac{-2}{3}\]

OpenStudy (asnaseer):

now we can use another log rule that states if:\[\log_a{b}=c\implies b=a^b\]

OpenStudy (asnaseer):

so, using this, we get:\[x=8^{\frac{-2}{3}}\] this can be simplified further

OpenStudy (asnaseer):

we can use this fact:\[a^{-b}=\frac{1}{a^ab\]to get:\[x=\frac{1}{8^{\frac{2}{3}}}\]

OpenStudy (asnaseer):

now, \(8=2^3\), so \(8^{\frac{1}{3}}=(2^3)^{\frac{1}{3}}=2\) so \(8^{\frac{2}{3}}=2^2=4\)

OpenStudy (asnaseer):

so finally, we get:\[x=\frac{1}{4}\]

OpenStudy (asnaseer):

do you understand the various steps above?

OpenStudy (anonymous):

a bit confusing

OpenStudy (asnaseer):

which step was confusing?

OpenStudy (asnaseer):

I can try and explain it in more detail

OpenStudy (anonymous):

I will try to do the other then I will let you know

OpenStudy (asnaseer):

ok - good luck

OpenStudy (anonymous):

I did this one take a look please: logx 16=5 X^5=16 x=\[\sqrt[5]{16}\]

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

you're a PRO now!

OpenStudy (asnaseer):

what about this one: log(x-3)+logx=log4 have you been able to do this yet?

OpenStudy (anonymous):

doing it now

OpenStudy (asnaseer):

keep in mind that if:\[\log(a)=\log(b)\implies a=b\]

OpenStudy (anonymous):

[(x-3)x]=4 (x-3)x=10^4 x^2-3x=10000 got scared by this big number am I on the right way?

OpenStudy (asnaseer):

:-)

OpenStudy (asnaseer):

not quite

OpenStudy (asnaseer):

you are correct up to: (x-3)x = 4

OpenStudy (asnaseer):

your next step after that was incorrect. where did you get the 10^ from?

OpenStudy (anonymous):

log of natural number

OpenStudy (asnaseer):

you should have ended up with:\[\log(x(x-3))=\log(4)\]which then leads to:\[x(x-3)=4\]

OpenStudy (asnaseer):

remember what I said above:\[\log(a)=\log(b)\implies a=b\]

OpenStudy (asnaseer):

if the log of 'a' equals the log of 'b', then 'a' must be equal to 'b'

OpenStudy (asnaseer):

do you understand this step?

OpenStudy (anonymous):

so is it going to be something like this: x^2-3x-4=0

OpenStudy (asnaseer):

exactly

OpenStudy (asnaseer):

this can be factorised

OpenStudy (asnaseer):

do you know to factorise?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

so what do you get for this equation?

OpenStudy (anonymous):

(x-4) (x+1)

OpenStudy (asnaseer):

perfect, so the solutions for 'x' are?

OpenStudy (anonymous):

x=4 or x=-1

OpenStudy (asnaseer):

correct - well done - you are a quick learner. many people would have given up by now. have faith in your abilities and you will be surprised by what you can achieve!

OpenStudy (anonymous):

thanks for you help

OpenStudy (asnaseer):

you're welcome.

OpenStudy (anonymous):

can you verify this one please: 3log x-1/2log y-log z logx^3-logy^1/2-logz logx^3-log\[\sqrt{y}-logz log x^3/z sqrt y